OpenStudy (anonymous):

when doing a power to a power i know you mutlipy the exponets but does the first nuber stay the same like (m^7)^2 i think its m^14 can someone check my answer?

OpenStudy (anonymous):

yes it stays the same

OpenStudy (anonymous):

\((b^n)^m=b^{mn}\) don't change the base

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[ (m^7)^2 = (\underbrace{mmmmmmm}_{7})^2 = \underbrace{mmmmmmmmmmmmmm}_{14} =m^{14} \]